# Forecasting the U.S. Renewable-Energy Mix with an ALR-BDARMA Compositional Time-Series Framework

### v2

@@ Abstract
Accurate forecasts of the U.S. renewable‑generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing‑market rules. We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025. The mean vector is modeled with a parsimonious VAR(2) in additive‑log‑ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
Forecast performance is assessed with a 61‑split rolling‑origin experiment that issues twelve‑month density forecasts from January 2019 to January 2024. Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift‑free ALR random walk—BDARMA lowers the mean continuous‑ranked probability score by 15–60 %, achieves component‑wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward. These results highlight BDARMA’s ability to deliver sharp, well‑calibrated probabilistic forecasts for multivariate renewable‑energy shares without sacrificing point precision.
Keywords: compositional time series; Dirichlet state‑space; Bayesian forecasting; renewable‑energy mix; seasonality.
@@ 1 Introduction
Electric‑sector decarbonisation hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve. Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S. Energy Information Administration, 2024; U.S. Energy Information Administration, 2025). Renewables already supply about one‑fifth of U.S. utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
@@ 2.1 Exploratory data analysis
Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD $=2.6$ pp), geothermal is quasi‑deterministic (SD $=0.17$ pp) and wood the most volatile component (SD $=5.1$ pp).
@@
To determine the minimum dynamic order in ALR space we fitted VAR(1) and VAR(2) models and applied Ljung–Box and Hosking portmanteau tests to the residuals (Table 2). A Ljung–Box residual diagnostic rejects the white‑noise null for coordinate $e_{3}$ under VAR(1) (p ¡ 0.001), whereas no coordinate is rejected under VAR(2) (smallest p‑value=.14). The residual ACF panels in Figure 4 show that the prominent spikes at lags 1–2 present under VAR(1) vanish when the second lag is added. At the system level, the portmanteau statistic at horizon 12 remains marginally significant; adding centered monthly dummies reduces $\chi^{2}$ from 628 to 431 ($p=0.006$). Because VAR(2) is the smallest specification to clear all short‑run autocorrelation and further lags inflate the parameter count without material gain, we adopt a VAR(2) mean and address any residual seasonality through exogenous Fourier terms.
@@
VAR(1) Ljung–Box $p$ 0.99 0.79 0.00 0.08 0.49 0.20
VAR(2) Ljung–Box $p$ 0.99 0.98 0.47 0.73 0.57 0.14
Portmanteau $\chi^{2}$ / $p$ 628/<0.001628\;/\;<0.001628 / < 0.001 (VAR(2))
@@ 3 Forecasting model
Let the monthly renewable‑energy mix be the $J\!=\!7$-component composition
@@
$\mathbf{y}_{t}\mid\boldsymbol{\mu}_{t},\varphi_{t}\;\sim\;\operatorname{Dirichlet}\bigl{(}\varphi_{t}\,\boldsymbol{\mu}_{t}\bigr{)},\qquad\boldsymbol{\mu}_{t}\in\mathcal{S}_{7},\;\varphi_{t}>0.$ (2)
@@
Each scalar element of $\mathbf{A}_{1},\mathbf{A}_{2},\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ receives an independent $\mathcal{N}(0,1)$ prior. Posterior inference proceeds via Hamiltonian Monte Carlo (four chains; 500 warm‑up and 500 retained iterations per chain) in Stan, yielding 2 00020002\,0002 000 draws that underpin all density‑forecast evaluations presented later.
@@ 3.1 Transform‑space VAR(2) (tVAR(2))
Parameters $(\mathbf{F}_{1},\mathbf{F}_{2},\boldsymbol{\delta},\boldsymbol{\Sigma})$ are estimated by ordinary least squares with the same seasonal regressors $\mathbf{X}_{t}$. Multi‑step forecasts are generated under the Gaussian innovation assumption and mapped back with ${\textnormal{alr}}^{-1}$.
@@ 4 Forecast–evaluation protocol
[…]
@@ 4.1 Generating predictive distributions
All four competitors are evaluated on Monte‑Carlo samples of equal size $M=2{,}000$ so that scoring rules are comparable.
@@
For every origin $s$ we retain the $M$ posterior draws $\{\boldsymbol{\theta}^{(m)}\}_{m=1}^{M}$ returned by the Hamiltonian Monte‑Carlo sampler. Each draw is propagated through the deterministic state equation (4) for $h=1{:}H$ steps, producing the latent mean $\boldsymbol{\mu}^{(m)}_{s,h}$; a single realisation $\mathbf{y}^{(m)}_{s,h}\sim\text{Dirichlet}(\phi^{(m)}_{s,h}\boldsymbol{\mu}^{(m)}_{s,h})$ is then generated from the observation density (LABEL:eq:obs). The empirical set $\mathcal{P}^{\text{BDARMA}}_{s,h}=\{\mathbf{y}^{(m)}_{s,h}\}_{m=1}^{M}$ constitutes the predictive distribution.
@@ 5.1 Forecast accuracy across horizons
Table 3 reports the mean continuous‑ranked probability score (CRPS) and Table 4 the mean Aitchison root‑mean‑square error (RMSE) for every horizon; the same information is visualised in Figures 6 and 7. When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field. Its mean CRPS at one step ahead is  0.00450.0045\,0.00450.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule. The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.

### v3

@@ Abstract
Accurate forecasts of the U.S. renewable-generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing-market rules. We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025. The mean vector is modeled with a parsimonious VAR(2) in additive-log-ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
Forecast performance is assessed with a 61-split rolling-origin experiment that issues twelve-month density forecasts from January 2019 to January 2024. Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift-free ALR random walk—BDARMA lowers the mean continuous-ranked probability score by 15–60 %, achieves component-wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward. These results highlight BDARMA’s ability to deliver sharp, well-calibrated probabilistic forecasts for multivariate renewable-energy shares without sacrificing point precision.
Keywords: compositional time series; Dirichlet state-space; Bayesian forecasting; renewable-energy mix; seasonality.
@@ 1 Introduction
Electric‑sector decarbonization hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve. Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S. Energy Information Administration, 2024; U.S. Energy Information Administration, 2025). Renewables already supply about one‑fifth of U.S. utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
@@ 2.1 Exploratory data analysis
Table 1 shows wide dispersion differences: hydro ranges from $7.5\%$ to $20.3\%$ (SD =2.6=2.6 pp), geothermal is quasi‑deterministic (SD =0.17=0.17 pp) and wood the most volatile component (SD =5.1=5.1 $p$).
@@
To determine the minimum dynamic order in ALR space we fitted VAR(1) and VAR(2) models and applied Ljung–Box and Hosking portmanteau tests to the residuals (Table 2). A Ljung–Box residual diagnostic rejects the white‑noise null for coordinate $e_{3}$ under VAR(1) (p < 0.001), whereas no coordinate is rejected under VAR(2) (smallest p‑value=.14). The residual ACF panels in Figure 4 show that the prominent spikes at lags 1–2 present under VAR(1) vanish when the second lag is added. At the system level, the portmanteau statistic at horizon 12 remains marginally significant; adding centered monthly dummies reduces $\chi^{2}$ from 628 to 431 ($p=0.006$). Because VAR(2) is the smallest specification to clear all short‑run autocorrelation and further lags inflate the parameter count without material gain, we adopt a VAR(2) mean and address any residual seasonality through exogenous Fourier terms.
@@
VAR(1) Ljung–Box pp 0.99 0.79 0.00 0.08 0.49 0.20
VAR(2) Ljung–Box pp 0.99 0.98 0.47 0.73 0.57 0.14
Portmanteau $\chi^{2}$ / pp 628/<0.001628\;/\;<0.001 (VAR(2))
@@ 3 Forecasting model
Let the monthly renewable‑energy mix be the J=7J\!=\!7-component composition
@@
$y_{t}\mid\mu_{t},\phi_{t}\sim\mathrm{Dirichlet}\!\big(\phi_{t}\mu_{t}\big),\quad\mu_{t}\in\mathcal{S}_{7},\;\phi_{t}>0.$ (2)
@@
Geometric preliminaries and evaluation mapping
Let $y_{t}\in\mathcal{S}_{7}$ denote the share vector and $e_{t}=\operatorname{alr}(y_{t})\in\mathbb{R}^{6}$ its additive log‑ratio (ALR) coordinates with biofuels as the reference part; $\operatorname{alr}^{-1}$ restores shares (see Eq. (1)). We model the mean in ALR space (Eq. (4)) and obtain predictive draws in share space from the Dirichlet observation layer (Eq. (2)). Forecasts are evaluated in two complementary spaces: CRPS in share space for joint sharpness and calibration (Eq. (5)), and clr‑based RMSE in Aitchison geometry for point accuracy (Eq. (6)). Because space choice induces different cross‑component dependencies, reporting both clarifies where improvements arise.
Reference‑free coordinates. The logistic‑normal (sometimes “ALN”) family places a Gaussian law on log‑ratio coordinates; the isometric log‑ratio (ILR) transform provides orthonormal, reference‑free coordinates with full metric equivalence on the simplex (Aitchison and Shen, 1980; Egozcue et al., 2003; Aitchison, 1986). This means that the DARMA data models with these three link functions are equivalent, provided the same transformation is applied to the priors. We retain ALR for interpretability and continuity with Eq. (1).
@@
Each scalar element of $\mathbf{A}_{1},\mathbf{A}_{2},\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ receives an independent $\mathcal{N}(0,1)$ prior. Posterior inference proceeds via Hamiltonian Monte Carlo (four chains; 500 warm‑up and 500 retained iterations per chain) in Stan, yielding 2 0002\,000 draws that underpin all density‑forecast evaluations presented later.
@@ 3.1 Transform‑space VAR(2) (tVAR(2))
Parameters $(\mathbf{F}_{1},\mathbf{F}_{2},\boldsymbol{\delta},\boldsymbol{\Sigma})$ are estimated by ordinary least squares with the same seasonal regressors $\mathbf{X}_{t}$. Multi‑step forecasts are generated under the Gaussian innovation assumption and mapped back with $\operatorname{alr}^{-1}$.
@@ 4 Forecast–evaluation protocol
Model comparison follows an expanding–window, rolling‑origin design that mirrors the workflow used by system operators and energy planners. Let $\tau_{s}$, $s=1,\dots,S$, denote the final observation included in estimation window $s$ and let $H$ denote the fixed forecast horizon ($H=12$). The first origin is $\tau_{1}=\text{2019‑01}$ and the last origin that still admits a twelve‑step look‑ahead is $\tau_{S}=\text{2024‑01}$, so $S=61$. At origin $s$ the estimation set is $\{\mathbf{y}_{t}:1\leq t\leq\tau_{s}\}$ while the verification set comprises $\{\mathbf{y}_{\,\tau_{s}+h}:h=1,\dots,H\}$.
@@ 4.1 Generating predictive distributions
All four competitors are evaluated on Monte‑Carlo samples of equal size M=2, [^0] M=2{,}000 so that scoring rules are comparable.
@@
For every origin $s$ we retain the $M$ posterior draws $\{\boldsymbol{\theta}^{(m)}\}_{m=1}^{M}$ returned by the Hamiltonian Monte‑Carlo sampler. Each draw is propagated through the deterministic state equation (4) for h=1:Hh=1{:}H steps, producing the latent mean $\boldsymbol{\mu}^{(m)}_{s,h}$; a single realization $\mathbf{y}^{(m)}_{s,h}\sim\text{Dirichlet}(\phi^{(m)}_{s,h}\boldsymbol{\mu}^{(m)}_{s,h})$ is then generated from the observation density (2). The empirical set $\mathcal{P}^{\text{BDARMA}}_{s,h}=\{\mathbf{y}^{(m)}_{s,h}\}_{m=1}^{M}$ constitutes the predictive distribution.
@@ 5.1 Forecast accuracy across horizons
Table 3 reports the mean continuous‑ranked probability score (CRPS) and Table 4 the mean Aitchison root‑mean‑square error (RMSE) for every horizon; the same information is visualised in Figures 6 and 7. When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field. Its mean CRPS at one step ahead is  0.0045\,0.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule. The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.